OpenStudy (anonymous):

I still don't understand why the solution for this diff equation is like this dy/(1-y)^2=1/(1-y) is there any explanation for this solution???thank you in advance

OpenStudy (anonymous):

\[\int\limits_{ }^{ }(\frac{ 1 }{ (1-y)^{2} })dy = \int\limits_{ }^{ }((1-y)^{-2})dy\] Just integrate it with a negative power, so when you increase the power by 1 it will go to -1 then divide by the new power so -(1-y)^-1 which equals -1/(1-y)

OpenStudy (anonymous):

thank you for your consideration, exactly we have the same integration result but I thin we got negative in the right side and for the task in the right side is possitive. would it be still the same meaning fo this sign?

OpenStudy (anonymous):

im only suggesting partial fractions..

OpenStudy (anonymous):

Using the substitution rule you should find that a -1 multiplication follows from the denominator (1-y).